# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ Appendix A ONLINE APPENDIX
Metrics
We consider metric of estimation error for the precision matrix as stated in Theorem 1. The model comparisons include several approaches from the literature. We consider the nonlinear shrinkage method of Ledoit and Wolf (2017)(NL-LW), POET method from Fan et al. (2013), and the nodewise (NW) method from Callot et al. (2021) We also tried single factor-nonlinear shrinkage method of Ledoit and Wolf (2017) but the results were worse than its no factor version of NL-LW so we decided to drop that. Our new method introduces several variations: RRE, PCR-3F, PCR-5F, PCR-20F and PCR-Adaptive. These denote the use of a RRE estimator, Principal component with fixed 3, 5, or 20 hidden factors, and an adaptive method to select the number of hidden factors, respectively. We do not include in the simulations, Caner et al. (2023) residual nodewise since it only uses observed factors. We also have done some very limited simulations with 3 factors with residual nodewise via the design here, treating unknown factors as known, and the results are similar to nodewise regression, hence omitted.
Results
Table OC.1- OC.2 shows the simulation results. The values in each cell show the average absolute estimation error for estimating the precision matrix. Each two-column block in the tables shows the result for a different sample size. The first column present results for $p=n/2$ (low-dimensional case), while the second corresponds to $p=3n/2$ (high-dimensional case). We ran 100 iterations in each simulation set up.
First, the tables show that our method achieves consistency as stated in Corollary 4 for adaptive PCR. Analyzing Table OC.1 with $K=3$, we observe that for $p=n/2$, with $n=100$, our Adaptive-PCR has an estimation error at 0.0583. This error declines to 0.0251 when $p=n/2$ and $n=400$. This indicates that as both $n$ and $p$ increase jointly, the estimation error decreases. Note that our RRE estimator in low dimensions (OLS) is consistent only when $p/n\to 0$ as shown in Corollary 5. However, it sill does well with $p=n/2$ in our Tables OC.1- OC.2. To give an example, from Table OC.1, the error is 0.4583 at $n=100$, and declines to 0.3369 at $n=400$. We also experimented with $p/n\to 0$ in our setup. As expected RRE error terms decline to zero fast, and these results are available from authors on demand.
Next, we consider which method achieves the smallest estimation errors. Our Adaptive-PCR dominates all other techniques except fixed PCR-K. In additional, our Adaptive-PCR effectively identifies the correct number of hidden factors in almost all of the scenarios. For instance, at $p=3n/2,n=400$ number of factor $K=5$ in Table OC.2, the estimation errors of both PCR-5F and Adaptive-PCR are the same at 0.0302, which also is the smallest estimation error among all methods.
Next, the number of hidden factors used in our PCR methods—whether underfitting or overfitting— impacts the estimation error. When the correct number of hidden factors is $K=3$, with $n=400$ and $p=3n/2$, the estimation error of PCR-3F is $0.0302$ in Table OC.1. However, overfitting the model by using $K=5,\,\text{or }20$ hidden factors increases the estimation error; for example, using PCR-20, the estimation error increases to $0.0401$. Underfitting, becomes problematic with $n=400$ in Table OC.2. In high dimensional scenario, PCR-5F has estimation error of 0.0302, whereas PCR-3F has error of 0.0524. As a side note to all of this, in RRE estimator, we see that estimation error for the low dimensional case is always higher than the double of high dimensional error scenario. This can be a symptom of double descent in RRE. To give an example, at Table OC.2 with $n=400$, the estimation error for RRE in low dimensional case is 0.3360, whereas in high dimensional case, RRE has error of 0.1417.
Table OC.1: Estimation Error,$K=3$
@@ Results
To examine the performance of our methods as the signal-to-noise ratio (SNR) changes, we analyze Figure 1 which shows the estimation error of the precision matrix in high-dimensional cases. We include the performance of RRE in blue and PCR-20F in red. From the figures, we observe that the performance of RRE and PCR improves as the SNR increases from 0 to 1. As the SNR increases beyond 1, the performance of PCR-20F continues to improve, while the performance of RRE stabilizes. Our PCR-20R consistently outperforms Inter-RRE across the entire SNR range for both low- and high-dimensional cases.
Figure 2: SNR and Estimation Errors: High Dimensions
@@
Here we provide two extra tables. One for asset size 100, and the other one is for 325.
Table OC.3: Monthly Portfolio Performance of 100 Stocks, $n_{I}=240$, $n-n_{I}=108$
@@
RNW-GIC-SF 0.0072 0.0460 0.1577 0.2934
RNW-GIC-3F 0.0089 0.0463 0.1912 0.3107
RNW-GIC-5F 0.0101 0.0491 0.2055 0.3604
NW-GIC 0.0097 0.0415 0.2338 0.1554
POET 0.0080 0.0462 0.1742 0.2339
Ledoit-Wolf 0.0092 0.0540 0.1711 0.5635
Inter-RRE 0.0065 0.0646 0.1009 1.1027
PCR-1F 0.0089 0.0395 0.2265 0.1632
PCR-3F 0.0113 0.0389 0.2900 0.3368
PCR-5F 0.0079 0.0424 0.1862 0.6514
PCR-7F 0.0126 0.0477 0.2649 0.8913
PCR-Adaptive 0.0067 0.0429 0.1561 0.5223
with transaction cost
RNW-GIC-SF 0.0076 0.0457 0.1654 NA
RNW-GIC-3F 0.0092 0.0461 0.1988 NA
RNW-GIC-5F 0.0105 0.0486 0.2167 NA
NW-GIC 0.0092 0.0415 0.2218 NA
POET 0.0083 0.0461 0.1801 NA
Ledoit-Wolf 0.0095 0.0536 0.1778 NA
Inter-RRE 0.0062 0.0645 0.0963 NA
PCR-1F 0.0089 0.0397 0.2244 NA
PCR-3F 0.0112 0.0390 0.2877 NA
PCR-5F 0.0081 0.0418 0.1939 NA
PCR-7F 0.0126 0.0471 0.2685 NA
PCR-Adaptive 0.0060 0.0431 0.1396 NA
Table OC.4: Monthly Portfolio Performance of 325 Stocks, $n_{I}=240$, $n-n_{I}=108$
@@
RNW-GIC-SF 0.0047 0.0442 0.1068 0.2982
RNW-GIC-3F 0.0060 0.0451 0.1324 0.3625
RNW-GIC-5F 0.0062 0.0474 0.1300 0.4650
NW-GIC 0.0095 0.0427 0.2228 0.1393
POET 0.0057 0.0438 0.1306 0.2105
Ledoit-Wolf 0.0023 0.0554 0.0415 0.8955
Inter-RRE 0.0108 0.0457 0.2356 0.3610
PCR-1F 0.0072 0.0373 0.1933 0.1391
PCR-3F 0.0066 0.0400 0.1660 0.3270
PCR-5F 0.0064 0.0535 0.1191 1.2584
PCR-7F 0.0096 0.0449 0.2146 2.0488
PCR-Adaptive 0.0084 0.0424 0.1983 0.1841
with transaction cost
RNW-GIC-SF 0.0050 0.0439 0.1144 NA
RNW-GIC-3F 0.0062 0.0449 0.1385 NA
RNW-GIC-5F 0.0064 0.0470 0.1370 NA
NW-GIC 0.0090 0.0427 0.2111 NA
POET 0.0059 0.0439 0.1349 NA
Ledoit-Wolf 0.0022 0.0552 0.0404 NA
Inter-RRE 0.0101 0.0458 0.2205 NA
PCR-1F 0.0071 0.0375 0.1898 NA
PCR-3F 0.0071 0.0393 0.1808 NA
PCR-5F 0.0048 0.0537 0.0898 NA
PCR-7F 0.0083 0.0438 0.1905 NA
PCR-Adaptive 0.0079 0.0425 0.1857 NA

### v2

@@ 6 Conclusion
Table C.1: Estimation Errors
K=3
@@
Here we provide the results of the empirics without transaction costs in Tables C.2-C.4. $SD$ is the standard deviation of the portfolio, and $Sharpe$ is the Sharpe ratio. Our methods still achieve the best Sharpe Ratio, turnover, standard deviation and return in Tables C.2-C.3 low dimension and moderate dimension cases. Table C.4, again our best method dominates the others in Return, Sharpe Ratio and Standard Deviation. In Turnover, nodewise has the best record. The first 6 rows represent the models in the literature, and the rest is our models. The bold faced numbers represent the winner in Sharpe Ratio.
Table C.2: Monthly Portfolio Performance of 10 Stocks, $_{I}=240, nn$$n_{I}=108$
@@
RNW-GIC-SF 0.0101 0.0526 0.1913 0.0953
RNW-GIC-3F 0.0101 0.0519 0.1945 0.0889
RNW-GIC-5F 0.0102 0.0510 0.2002 0.0876
NW-GIC 0.0106 0.0505 0.2093 0.0888
POET 0.0104 0.0507 0.2051 0.0700
Ledoit-Wolf 0.0100 0.0522 0.1914 0.0927
Inter-RRE 0.0104 0.0530 0.1957 0.0929
PCR-1F 0.0109 0.0490 0.2221 0.0518
PCR-3F 0.0105 0.0492 0.2142 0.0576
PCR-5F 0.0106 0.0501 0.2125 0.0948
PCR-7F 0.0105 0.0507 0.2075 0.0795
PCR-Adaptive 0.0108 0.0489 0.2211 0.0514
Table C.3: Monthly Portfolio Performance of 200 Stocks, $_{I}=240, nn$$n_{I}=108$
@@
RNW-GIC-SF 0.0053 0.0442 0.1205 0.2761
RNW-GIC-3F 0.0067 0.0447 0.1496 0.3171
RNW-GIC-5F 0.0072 0.0484 0.1483 0.3996
NW-GIC 0.0093 0.0429 0.2169 0.1517
POET 0.0057 0.0430 0.1336 0.1989
Ledoit-Wolf 0.0048 0.0521 0.0928 0.7320
Inter-RRE -0.0018 0.1168 -0.0158 6.1638
PCR-1F 0.0075 0.0376 0.2002 0.1203
PCR-3F 0.0082 0.0382 0.2151 0.2860
PCR-5F 0.0081 0.0557 0.1462 1.0298
PCR-7F 0.0128 0.0433 0.2964 1.0439
PCR-Adaptive 0.0081 0.0435 0.1865 0.4231
Table C.4: Monthly Portfolio Performance of 357 Stocks, $_{I}=240, nn$$n_{I}=108$
Estimator Maximum Sharpe Ratio
Return SD Sharpe Turnover
without transaction cost
RNW-GIC-SF 0.0046 0.0432 0.1068 0.2857
RNW-GIC-3F 0.0057 0.0440 0.1297 0.3580
RNW-GIC-5F 0.0056 0.0461 0.1209 0.4507
NW-GIC 0.0093 0.0422 0.2202 0.1333
POET 0.0058 0.0424 0.1364 0.2009
Ledoit-Wolf 0.0015 0.0534 0.0280 0.9073
Inter-RRE 0.0106 0.0453 0.2338 0.2725
PCR-1F 0.0073 0.0364 0.2004 0.1349
PCR-3F 0.0054 0.0366 0.1470 0.2798
PCR-5F 0.0056 0.0528 0.1056 1.6010
PCR-7F 0.0095 0.0506 0.1870 1.8931
PCR-Adaptive 0.0085 0.0420 0.2020 0.2608
Discussion of Literature
Several studies in the statistics literature have studied precision matrix estimation. An excellent survey is by Cai et al. (2016).
In Section 3 of their paper, they provide two major approaches to precision matrix estimation. The first approach is neighborhood-based and closely related to nodewise regression, where a lasso regression is performed for each variable against all others. Alternatives to lasso, such as the Dantzig selector, can also be used within this framework (see Cai and Liu (2011), Callot et al. (2021)). The second major approach is based on penalized likelihood estimation. As shown by Bertsimas et al. (2020), lasso-type penalties in this context can induce shrinkage bias and lead to overly sparse precision matrix estimates, with many zero entries in each row. In a penalized likelihood based setting, they reformulate the problem as a mixed-integer optimization problem, which is solvable using convex optimization techniques. This approach produces better results than lasso based approaches. Sparsity of precision matrix is used in these two main approaches. A distinct approach, commonly used in finance and statistics, derives the precision matrix by inverting the covariance matrix using the Sherman–Morrison–Woodbury (SMW) formula. In this framework, sparsity is assumed in the covariance matrix of the idiosyncratic errors, while asset returns are modeled through a factor structure (see Fan et al. (2013), Caner et al. (2023)). Our approach is different from the approaches discussed in the literature. First of all, we are not assuming exact sparsity in the precision matrix of outcomes. We are also not using SMW formula to invert the covariance matrix.
This allows us to bypass imposing a structure on the outcomes utilized in the covariance inversion. Instead we use a new precision matrix formula without sparsity and rely on hidden factors to generate each outcome. This effectively ensures that we do not benefit from the SMW matrix inversion formula. To be clear, we use model estimation-free dense estimation, but consistency in high dimensions is achieved through hidden factor based dimension reduction without a need to estimate factors or identify them.
There are also other technical and conceptual differences that are worth highlighting. First, our proposed class of estimators is general and in addition to the PCR based ones it also includes the ridgeless regression estimator (RRE). Second, our signal assumption is much weaker as we do not need the minimum signal to grow at the same rate as the number of assets in the portfolio. Moreover, as the root of our approach resembles a finite sample analysis, it also allows to explicitly tie all the estimation errors to signal-to-noise ratio thereby making its importance in the rate of convergence estimates clearly visible.
The precision matrix estimation is different than the risk analysis of a single outcome-asset. We also provide an explicit connection in our estimation error upper bounds to signal-to-noise ratio. The results in subcases are different also in a risk analysis versus precision matrix estimation. Bing et al. (2021), Bunea et al. (2022) show that in certain very high dimensional cases, ridgeless regression (RRE) can perform better than adaptive PCR in terms of risk asymptotically. However, in case of estimation errors of rows of precision matrix errors (in $l_{2}$ norm) we show adaptive PCR can perform better than RRE even in certain low-moderate dimension cases. These are due to different proofs for risk versus precision matrix estimation.
Hidden Factors
The reason we use factor models is for dimension reduction. First of all, we show that precision matrix estimation is related to linear regression and that can be related to hidden factors. Each economic or financial variable can be related to latent factors and this can be thought as confounding variables as in causality related literature such as in Kaltenpoth and Vreeke (2023), Angrist and Pischke (2009), Imbens and Rubin (2015). Also the true variables of interest may be measured with errors, and latent factors can represent the ”true” variables rather than the reported ones, especially in emerging markets. Variable such as culture, sentiment, market can be extremely valuable in certain macro, financial data sets. These hidden factors in a linear model can be used to explain international stock returns such as in Singapore, Malaysia, USA as seen in Brooks and Del Negro (2004).
Another key aspect of using hidden factors is their effectiveness as a tool for dimension reduction in high-dimensional econometrics. This stems from the fact that the number of observed outcomes typically exceeds the number of underlying hidden factors. In our paper, we allow both the number of outcomes and the number of hidden factors to grow with the sample size.